OpenStudy (anonymous):

solve 5p-4+3p=2+7p+3

OpenStudy (angela210793):

5p-4+3p-2-7p+3=0 -2p=3 p=-3/2=-1.5

OpenStudy (anonymous):

p = 9

OpenStudy (angela210793):

huh? how 9?

OpenStudy (angela210793):

Oops.....yes ur right so sorry